-2(7-a) I Suppsose This Is easy But I thought it out and am asking, would the answer be -14-a?
Answer by stanbon(75887)

-2(7-a) I Suppose This Is easy But I thought it out and am asking, would the answer be -14-a?
-----
No::
-2(7-a) = -2*7 + -2*-a = -14 + 2a
--------------------
Cheers,
Stan H.
------------